OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! f is a differentiable function for all x. Which of the following statements must be true? I. the derivative with respect to x of the integral from 0 to 4 of f of x, dx is equal to f of x II. d/dx of integral from x to 4 of f of t d t equals negative f of x III. the integral from 4 to x of f prime of x, dx is equal to negative f of x a. II only b. III only c. I and II only d. II and III only

OpenStudy (aryana_maria2323):

Can you please help? @agent0smith @ganeshie8 @Kainui @mathmale @MrNood @Nnesha @zepdrix

OpenStudy (xguardians):

I is correct a believe, due to the fundamental theory of calculus.

OpenStudy (xguardians):

III is wrong.

OpenStudy (aryana_maria2323):

Okay so then my answer is C? Since it is the only one with "I" in it?

OpenStudy (xguardians):

It's either A or C I believe.

OpenStudy (xguardians):

Not sure why f(t) is in there...

OpenStudy (peachpi):

I don't think I is true. The integral from 0 to 4 is a constant, so its derivative is 0

OpenStudy (aryana_maria2323):

Can you please help me with the one more question? @xGuardians

OpenStudy (xguardians):

@peachpi That's incorrect. Why would it be a constant?

OpenStudy (xguardians):

@aryana_maria2323 Yes.

OpenStudy (peachpi):

FTC, part I: \[\frac{ d }{ dt }\int\limits_{a}^{x}f(t)dt=f(x)\] I isn't true because the limits on the integral are both constants II is true because if you reverse the order of the integral's limits, the result has the opposite sign III, I'm not entirely sure about

OpenStudy (aryana_maria2323):

What are the limits of integration if the summation the limit as n goes to infinity of the summation from k equals 1 to n of the product of the quantity of the square of 2 plus 7 times k over n and the quotient of 7 and n is written as a definite integral with integrand x2? a. 0,2 b. 1,7 c. 2,9 d. 2,7

OpenStudy (peachpi):

@xGuardians it's constant because that's a definite integral, which will return a specific number, not a function

OpenStudy (peachpi):

On the previous question, I'd lean toward III being false because there's nothing in that integral that would make it negative.

OpenStudy (aryana_maria2323):

Can you please help me with my new question?

OpenStudy (peachpi):

can you write that out with the equation tool?

OpenStudy (aryana_maria2323):

I did?

OpenStudy (peachpi):

if the summation the limit as n goes to infinity of the summation from k equals 1 to n of the product of the quantity of the square of 2 plus 7 times k over n and the quotient of 7 and n is written as a definite integral with integrand x2? ^that's what I'm seeing. Only text

OpenStudy (aryana_maria2323):

Okay one second

OpenStudy (aryana_maria2323):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n} [2+\frac{ 7k }{ n }]^2 (\frac{ 7 }{ n })\]

OpenStudy (peachpi):

sorry, I don't know how to do that